# Test-Time Alignment of LLMs via Sampling-Based Optimal Control in pre-logit space

### v1

@@ Abstract
Test-time alignment of large language models (LLMs) attracts attention because fine-tuning LLMs requires high computational costs. In this paper, we propose a new test-time alignment method called adaptive importance sampling on pre-logits (AISP) on the basis of the sampling-based model predictive control with the stochastic control input. AISP applies the Gaussian perturbation into pre-logits, which are outputs of the penultimate layer, so as to maximize expected rewards with respect to the mean of the perturbation. We demonstrate that the optimal mean is obtained by importance sampling with sampled rewards. AISP outperforms best-of-n sampling in terms of rewards over the number of used samples and achieves higher rewards than other reward-based test-time alignment methods.
@@ 1 Introduction
In this paper, we propose a new test-time alignment inspired by sampling-based control methods without a training process. Traditional optimal control theory can optimize input trajectories without any training process by solving differential equations such as the Pontryagin’s maximum principle. However, these methods are not applicable to LLM alignment because LLMs are nonlinear, complicated and large-scale systems (Chen et al., 2024). For such systems, sampling-based model predictive control has been advanced by leveraging the parallel computing capabilities of GPUs (Williams et al., 2018; 2017). Therefore, we adopt this optimal control technique in the LLM alignment by incorporating adaptive importance sampling. First, we formalize LLM alignment as a stochastic control problem where the control input is a stochastic perturbation on pre-logits, which are outputs of the penultimate layer of the LLM. In our formulation, the perturbation follows a Gaussian distribution. According to this, the distribution of pre-logit sequences also becomes a Gaussian distribution and can be written by the closed form, unlike token sequence distributions. Next, we derive the optimal distribution through the free energy that bounds this problem. Since this distribution is intractable, we approximate it by using importance sampling where the weighting function can be easily computed thanks to the Gaussian assumption. We iteratively update the proposal distribution via adaptive importance sampling (Kloek & Van Dijk, 1978; Cappé et al., 2004; Bugallo et al., 2017) because naïve importance sampling can require a large number of effective samples due to the vast pre-logit sequence space. Therefore, our method is called adaptive importance sampling on pre-logits (AISP, Fig. 1). After explanation of AISP, we discuss the connection between the Gaussian assumption and the last softmax layer in neural networks. Additionally, we reveal that AISP becomes equivalent to BoN with the specific sampling strategy in the limit of a hyperparameter. Experiments demonstrate that AISP increases reward values faster than BoN in terms of the number of used generated samples. Additionally, AISP also outperforms RE-Control even though it does not require training dataset collection in advance. Since AISP requires fewer samples than BoN, we also evaluate Batched AISP, which simultaneously handles multiple prompts with small samples, and confirm that Batched AISP can outperform BoN under the same iterations.
@@ 2.1 Best-of-N sampling
$\textstyle\bm{\mathrm{y}}_{\mathrm{BoN}}=\operatorname*{arg\,max}_{\bm{\mathrm{y}}\in\mathcal{Y}_{N}}r(\bm{\mathrm{x}},\bm{\mathrm{y}}).$ (1)
@@
$\textstyle\max_{\pi(\cdot|\bm{\mathrm{x}})}\mathbb{E}_{\bm{\mathrm{y}}\sim\pi(\cdot|\bm{\mathrm{x}})}r(\bm{\mathrm{x}},\bm{\mathrm{y}})-\lambda D_{KL}(\pi(\cdot|\bm{\mathrm{x}})|P_{\mathrm{LLM}}(\cdot|\bm{\mathrm{x}})).$ (2)
$D_{KL}(\pi(\cdot|\bm{\mathrm{x}})|P_{\mathrm{LLM}}(\cdot|\bm{\mathrm{x}}))$ prevents $\pi(\cdot|\bm{\mathrm{x}})$ from moving far away from the base LLM $P_{\mathrm{LLM}}(\cdot|\bm{\mathrm{x}})$. Equation (2) has a closed solution (Beirami et al., 2024; Korbak et al., 2022; Go et al., 2023):
@@
where $\eta\!=\!\sum_{\bm{\mathrm{y}}}P_{\mathrm{LLM}}(\bm{\mathrm{y}}|\bm{\mathrm{x}})\mathrm{exp}(\frac{1}{\lambda}r(\bm{\mathrm{x}},\bm{\mathrm{y}}))$ is a normalization constant, which is hard to estimate (Rafailov et al., 2023).
@@ 2.2 RE-Control
where $\bm{W}_{\mathrm{LLM}}$ and $\bm{b}_{\mathrm{LLM}}$ are the parameter of the last linear layer of the LLM. $\bm{z}_{t}\in\mathbb{R}^{d}$ is called pre-logit [^1] Though Kong et al. (2024) called $\bm{z}_{t}$ logit, we call it pre-logit to distinguish it from the the input to the softmax function: $\bm{W}_{\mathrm{LLM}}\bm{z}_{t}+\bm{b}_{\mathrm{LLM}}$., which is the output vector of the penultimate layer of the LLM: $\bm{z}_{t}=\phi_{\mathrm{LLM}}(\bm{y}_{<t})$ where $\bm{y}_{<t}=[y_{0},\dots,y_{t-1}]$ are the past tokens including the input prompt $\mathbf{x}$. $\phi_{\mathrm{LLM}}(\cdot)$ contains an embedding layer and attention layers. In this formulation, $\bm{u}_{t}$ is optimized through the gradient ascent to maximize the value function $V(\bm{z}_{t})$, which evaluates the current state of LLMs to maximize rewards at the terminal. However, this value function needs to be trained on the dataset, which are composed of various states, responses, and rewards: $D_{V}\!=\!\{(\bm{z}_{0:T}^{i},\mathbf{y}^{i},r(\mathbf{x}^{i},\mathbf{y}^{i}))\}_{i=1}^{M}$. In fact, Kong et al. (2024) uses 349,000 training prompts in SHP (Ethayarajh et al., 2022) to collect them, which incurs storage costs and training time. Instead of training the value function, we adopt a sampling-based optimal control to LLM alignment.
@@ 3 Proposed method: AISP
To maximize rewards, we consider applying the control theory to LLM alignment similar to Kong et al. (2024), but without training. In fact, traditional optimal control methods do not require any training process because the optimal input trajectories are derived by solving differential equations such as the Pontryagin’s maximum principle. However, such methods are ineffective for LLM alignment because LLMs are nonlinear large-scale systems (Chen et al., 2024). For such systems, recent optimal control methods have incorporated a sampling-based approach with model predictive control by considering stochastic input. Thus, we adopt the stochastic optimal control method called model predictive path integral control (MPPI) (Williams et al., 2018; 2017) to LLM alignment. First, we formalize our problem and explain the closed solution. Since it is an intractable distribution, we present adaptive importance sampling to solve this problem. Next, we discuss the connection between the assumption in pre-logit and softmax function, and the connection with BoN. Finally, we explain the details of implementation.
@@ 3.1 Problem Formulation
Whereas RE-Control (Eq. (4)) uses the deterministic input $\bm{u}_{t}$, we apply the stochastic control input $\bm{v}_{t}$ with mean $\bm{u}_{t}$ to LLMs and optimize $\bm{u}_{t}$. Specifically, we inject a Gaussian noise $\bm{v}_{t}\!\in\!\mathbb{R}^{d}$ to pre-logit $\bm{z}_{t}\!=\!\phi_{\mathrm{LLM}}(\bm{y}_{<t})$ for the time interval $t\!\in\![1,\tau]$. The input prompt $\mathbf{x}$ corresponds to $\bm{y}_{<1}$. Then, a pre-logit follows a Gaussian distribution, and the $t$-th token is given by
$y_{t}=\begin{cases}\mathrm{softmax}(\bm{W}_{\mathrm{LLM}}(\bm{z}_{t}+\bm{v}_{t})+\bm{b}_{\mathrm{LLM}}),\bm{v}_{t}\sim\mathcal{N}(\bm{u}_{t},\sigma^{2}\bm{I}),&\mathrm{for}~~1\leq t\leq\tau,\\ \mathrm{softmax}(\bm{W}_{\mathrm{LLM}}\bm{z}_{t}+\bm{b}_{\mathrm{LLM}}),&\mathrm{for}~~\tau<t.\end{cases}$ (5)
where $\sigma^{2}\bm{I}$ is a covariance matrix with a fixed variance $\sigma^{2}\!\in\!\mathbb{R}$. This can be interpreted as the distribution of the pre-logit $\hat{\bm{z}}_{t}$ in AISP is given by $p(\hat{\bm{z}}_{t}|\bm{y}_{<t})\!=\!\mathcal{N}(\phi_{\mathrm{LLM}}(\bm{y}_{<t})+\bm{u}_{t},\sigma^{2}\bm{I})$ for $t\!\in\![1,\tau]$. The distribution of the input trajectory $V\!=\![\bm{v}_{1},\dots,\bm{v}_{\tau}]\!\in\!\mathbb{R}^{d\times\tau}$ is a joint Gaussian distribution:
$\textstyle q(V|U,\sigma^{2})=\frac{1}{(2\pi\sigma^{2})^{\frac{d\tau}{2}}}\mathrm{exp}\left(-\frac{1}{2\sigma^{2}}\sum_{t=1}^{\tau}(\bm{v}_{t}-\bm{u}_{t})^{\top}(\bm{v}_{t}-\bm{u}_{t})\right),$ (6)
@@
$\min_{U}J(\bm{\mathrm{x}},U)=\min_{U}-\mathbb{E}_{V\sim\mathbb{Q}_{U,\sigma^{2}}}\left[r(\bm{\mathrm{x}},\bm{\mathrm{y}}(V))\right]+\lambda\mathbb{D}_{\mathrm{KL}}(\mathbb{Q}_{U,\sigma^{2}}|\mathbb{P})$ (7)
@@
$\textstyle p(V|\bm{0},\sigma^{2})=\frac{1}{(2\pi\sigma^{2})^{\frac{d\tau}{2}}}\mathrm{exp}\left(-\frac{1}{2\sigma^{2}}\sum_{t=1}^{\tau}\bm{z}_{t}^{\top}\bm{z}_{t}\right).$ (8)
This implies that we assume the pre-logit distribution of the base LLM following $p_{\mathrm{LLM}}(\bm{z}_{t}|\bm{y}_{<t})\!=\!\mathcal{N}(\phi_{\mathrm{LLM}}(\bm{y}_{<t}),\sigma^{2}\bm{I})$. The KL-divergence $\mathbb{D}_{\mathrm{KL}}(\mathbb{Q}_{U,\sigma^{2}}|\mathbb{P})$ is given by $\mathbb{D}_{\mathrm{KL}}(\mathbb{Q}_{U,\sigma^{2}}|\mathbb{P})\!=\!1/2\sigma^{2}\sum_{t=1}^{\tau}\!\bm{u}^{\top}_{t}\bm{u}_{t}$. After the optimization, we can obtain the optimal reward $\bm{\mathrm{y}}_{\mathrm{AISP}}$ as $\bm{\mathrm{y}}_{\mathrm{AISP}}\!=\!\bm{\mathrm{y}}(V\!=\!U^{*})$ where $U^{*}$ is the solution of Eq. (7). Since we can generate several candidate responses at the last, we select $\bm{\mathrm{y}}_{\mathrm{AISP}}\!=\!\mathrm{arg}\!\max_{V\in\mathcal{V}}\bm{\mathrm{y}}(V)$ where $\mathcal{V}\!=\!\{V^{i}|V^{i}\!\sim\!q(V|U^{*},\sigma^{2})\}$ instead of just using $\bm{\mathrm{y}}(V\!=\!U^{*})$.
@@ 3.2 Free energy and Optimal distribution
Optimization problems such as Eq. (7) correspond to the optimal control problems called model predictive path integral control (MPPI) (Williams et al., 2018; 2017) and can be solved by considering a certain free energy. To optimize Eq. (7), we consider the following free energy:
@@
By using $\mathbb{Q}$ and Jensen’s inequality, we have the following result:
@@
All of the proofs can be found in  A. This theorem shows that the free energy Eq. (9) is the lower bound of the objective function Eq. (7), and the optimal density function is given by Eq. (10). The optimal density function Eq. (10) is intractable, and it is difficult to obtain directly. Next, we show how to approximate it by using importance sampling. Note that while the similar result Eq. (3) has been presented for the distribution of responses, Eq. (10) is related to the distribution of pre-logits. This difference makes the optimization easy because $p(V)=p(V|\bm{0},\sigma)$ in Eq. (10) is a tractable distribution, and its sampling is easy as written in the next subsection.
@@ 3.3 Adaptive importance sampling
We consider to approximate $\mathbb{Q}^{*}$ of Eq. (10) by the Gaussian distribution of Eq. (6) though importance sampling (Robert et al., 1999; Kloek & Van Dijk, 1978; Cappé et al., 2004; Bugallo et al., 2017). We recall the following theorem, which was established in Williams et al. (2018):
@@
$\textstyle w(V)$ $\textstyle=\frac{1}{\eta}\mathrm{exp}\left(\frac{1}{\lambda}r(\bm{\mathrm{x}},\bm{\mathrm{y}}(V))-\frac{1}{\sigma^{2}}\sum_{t=1}^{\tau}\hat{\bm{u}}_{t}^{\top}\bm{v}_{t}-\frac{1}{2}\hat{\bm{u}}_{t}^{\top}\hat{\bm{u}}_{t}\right).$ (12)
@@
Note that our formulation is followed by MPPI (Williams et al., 2018) and can actually be extended to MPPI on pre-logits by changing sampling methods. While MPPI moves prediction and control windows by determining and applying $\hat{\bm{u}}_{1}$ to the control target for each iteration, AISP uses a fixed control window $t\!\in\![0,\tau]$. This is because once moving windows fix the prefix tokens, generated responses lose diversity. AISP can explore large response spaces by using the fixed window starting t=0t\!=\!0 and adaptive importance sampling.

### v2

@@ Abstract
Test-time alignment of large language models (LLMs) attracts attention because fine-tuning of LLMs requires high computational costs. In this paper, we propose a new test-time alignment method called adaptive importance sampling on pre-logits (AISP) on the basis of the sampling-based model predictive control with the stochastic control input. AISP applies the Gaussian perturbation into pre-logits, which are outputs of the penultimate layer, so as to maximize expected rewards with respect to the mean of the perturbation. We demonstrate that the optimal mean is obtained by importance sampling with sampled rewards. AISP outperforms best-of-n sampling in terms of rewards over the number of used samples and achieves higher rewards than other reward-based test-time alignment methods.
LLM, Model predictive control
@@ 1 Introduction
In this paper, we propose a new test-time alignment inspired by sampling-based control methods without a training process. Traditional optimal control theory can optimize input trajectories without any training process by solving differential equations such as the Pontryagin’s maximum principle. However, these methods are not applicable to LLM alignment because LLMs are nonlinear, complicated and large-scale systems (Chen et al., 2024). For such systems, sampling-based model predictive control has been advanced by leveraging the parallel computing capabilities of GPUs (Williams et al., 2018, 2017). Therefore, we adopt this optimal control technique in the LLM alignment by incorporating adaptive importance sampling. First, we formalize LLM alignment as a stochastic control problem where the control input is a stochastic perturbation on pre-logits, which are outputs of the penultimate layer of the LLM. In our formulation, the perturbation follows a Gaussian distribution. According to this, the distribution of pre-logit sequences also becomes a Gaussian distribution and can be written by the closed form, unlike token sequence distributions. Next, we derive the optimal distribution through the free energy that bounds this problem. Since this distribution is intractable, we approximate it by using importance sampling where the weighting function can be easily computed thanks to the Gaussian assumption. We iteratively update the proposal distribution via adaptive importance sampling (Kloek and Van Dijk, 1978; Cappé et al., 2004; Bugallo et al., 2017) because naïve importance sampling can require a large number of effective samples due to the vast pre-logit sequence space. Therefore, our method is called adaptive importance sampling on pre-logits (AISP, Fig. 1). After explanation of AISP, we discuss the connection between the Gaussian assumption and the last softmax layer in neural networks. Additionally, we reveal that AISP becomes equivalent to BoN with the specific sampling strategy in the limit of a hyperparameter. Experiments demonstrate that AISP increases reward values faster than BoN in terms of the number of used generated samples. Additionally, AISP also outperforms RE-Control even though it does not require training dataset collection in advance. Since AISP requires fewer samples than BoN, we also evaluate Batched AISP, which simultaneously handles multiple prompts with small samples, and confirm that Batched AISP can outperform BoN under the same iterations.
@@ 2.1 Best-of-N sampling
$\bm{\mathrm{y}}_{\mathrm{BoN}}=\operatorname*{arg\,max}_{\bm{\mathrm{y}}\in\mathcal{Y}_{N}}r(\bm{\mathrm{x}},\bm{\mathrm{y}}).$ (1)
@@
$\max_{\pi(\cdot|\bm{\mathrm{x}})}\mathbb{E}_{\bm{\mathrm{y}}\sim\pi(\cdot|\bm{\mathrm{x}})}r(\bm{\mathrm{x}},\bm{\mathrm{y}})-\lambda D_{KL}(\pi(\cdot|\bm{\mathrm{x}})|P_{\mathrm{LLM}}(\cdot|\bm{\mathrm{x}})).$ (2)
The KL divergence $D_{KL}(\pi(\cdot|\bm{\mathrm{x}})|P_{\mathrm{LLM}}(\cdot|\bm{\mathrm{x}}))$ prevents $\pi(\cdot|\bm{\mathrm{x}})$ from moving far away from the base LLM $P_{\mathrm{LLM}}(\cdot|\bm{\mathrm{x}})$. Equation (2) has a closed solution (Beirami et al., 2024; Korbak et al., 2022; Go et al., 2023):
@@
where $\eta\!=\!\sum_{\bm{\mathrm{y}}}P_{\mathrm{LLM}}(\bm{\mathrm{y}}|\bm{\mathrm{x}})\mathrm{exp}(\frac{1}{\lambda}r(\bm{\mathrm{x}},\bm{\mathrm{y}}))$. $\eta$ is a normalization constant, which is hard to estimate (Rafailov et al., 2023).
@@ 2.2 RE-Control
where $\bm{W}_{\mathrm{LLM}}$ and $\bm{b}_{\mathrm{LLM}}$ are the parameter of the last linear layer of the LLM. $\bm{z}_{t}\in\mathbb{R}^{d}$ is called pre-logit [^1] Though Kong et al. (2024) called $\bm{z}_{t}$ logit, we call it pre-logit to distinguish it from the the input to the softmax function: $\bm{W}_{\mathrm{LLM}}\bm{z}_{t}+\bm{b}_{\mathrm{LLM}}$., which is the output vector of the penultimate layer of the LLM: $\bm{z}_{t}\!=\!\phi_{\mathrm{LLM}}(\bm{y}_{<t})$ where $\bm{y}_{<t}\!=\![y_{0},\dots,y_{t-1}]$ is a past token sequence including the input prompt $\mathbf{x}$. $\phi_{\mathrm{LLM}}(\cdot)$ contains an embedding layer and attention layers. In this formulation, $\bm{u}_{t}$ is optimized through the gradient ascent to maximize the value function $V(\bm{z}_{t})$, which evaluates the current state of LLMs to maximize rewards at the terminal. However, this value function needs to be trained on the dataset, which are composed of various states, responses, and rewards: $D_{V}\!=\!\{(\bm{z}_{0:T}^{i},\mathbf{y}^{i},r(\mathbf{x}^{i},\mathbf{y}^{i}))\}_{i=1}^{M}$. In fact, Kong et al. (2024) uses 349,000 training prompts in SHP (Ethayarajh et al., 2022) to collect them, which incurs storage costs and training time. Instead of training the value function, we adopt a sampling-based optimal control to LLM alignment.
@@ 3 Proposed method: AISP
To maximize rewards, we consider applying the control theory to LLM alignment similar to (Kong et al., 2024), but without training. In fact, traditional optimal control methods do not require any training process because the optimal input trajectories are derived by solving differential equations such as the Pontryagin’s maximum principle. However, such methods are ineffective for LLM alignment because LLMs are nonlinear large-scale systems (Chen et al., 2024). For such systems, recent optimal control methods have incorporated a sampling-based approach with model predictive control by considering stochastic input. Thus, we adopt the stochastic optimal control method called model predictive path integral control (MPPI) (Williams et al., 2018, 2017) to LLM alignment. First, we formalize our problem and explain the closed solution. Since it is an intractable distribution, we present adaptive importance sampling to solve this problem. Next, we discuss the connection between the assumption in pre-logit and softmax function, and the connection with BoN. Finally, we explain the details of implementation.
@@ 3.1 Problem Formulation
Whereas RE-Control (Eq. (4)) uses the deterministic input $\bm{u}_{t}$, we apply the stochastic control input $\bm{v}_{t}$ with mean $\bm{u}_{t}$ to LLMs and optimize $\bm{u}_{t}$. Specifically, we inject a Gaussian noise $\bm{v}_{t}\!\in\!\mathbb{R}^{d}$ to pre-logit $\bm{z}_{t}\!=\!\phi_{\mathrm{LLM}}(\bm{y}_{<t})$ for the time interval $t\!\in\![1,\tau]$, where $\tau$ is practically set to maximum generation length. The input prompt $\mathbf{x}$ corresponds to $\bm{y}_{<1}$. Then, the $t$-th token is given by
$\!y_{t}$ $\!=\!\operatorname*{arg\,max}_{i}\!\left[\mathrm{softmax}(\bm{W}_{\mathrm{LLM}}(\bm{z}_{t}+\bm{v}_{t})+\bm{b}_{\mathrm{LLM}})\right]_{i},$ (5)
where $\bm{v}_{t}\!\sim\!\mathcal{N}(\bm{u}_{t},\sigma^{2}\bm{I})$ and $\sigma^{2}\bm{I}$ is a covariance matrix with a fixed variance $\sigma^{2}\!\in\!\mathbb{R}$. This can be interpreted as the distribution of the pre-logit $\hat{\bm{z}}_{t}$ is given by $p(\hat{\bm{z}}_{t}|\bm{y}_{<t})\!=\!\mathcal{N}(\phi_{\mathrm{LLM}}(\bm{y}_{<t})+\bm{u}_{t},\sigma^{2}\bm{I})$ for $t\!\in\![1,\tau]$. The distribution of the input trajectory $V\!=\![\bm{v}_{1},\dots,\bm{v}_{\tau}]\!\in\!\mathbb{R}^{d\times\tau}$ is a joint Gaussian distribution:
$\textstyle q(V|U,\!\sigma^{2})\!=\!\frac{1}{(2\pi\sigma^{2})^{\frac{d\tau}{2}}}\mathrm{exp}\!\left(\!-\frac{\sum_{t=1}^{\tau}(\bm{v}_{t}-\bm{u}_{t})^{\top}(\bm{v}_{t}-\bm{u}_{t})}{2\sigma^{2}}\!\right)\!,$ (6)
@@
$\min_{U}J(\bm{\mathrm{x}},U)$
$=$ $\min_{U}-\mathbb{E}_{V\sim\mathbb{Q}_{U,\sigma^{2}}}\!$ $\left[r(\bm{\mathrm{x}},\bm{\mathrm{y}}(V))\right]+\lambda\mathbb{D}_{\mathrm{KL}}(\mathbb{Q}_{U,\sigma^{2}}|\mathbb{P})$ (7)
@@
$\textstyle p(V|\bm{0},\sigma^{2})=\frac{1}{(2\pi\sigma^{2})^{\frac{d\tau}{2}}}\mathrm{exp}\left(-\frac{1}{2\sigma^{2}}\sum_{t=1}^{\tau}\bm{v}_{t}^{\top}\bm{v}_{t}\right).$ (8)
This implies that we assume the pre-logit distribution of the base LLM following $p_{\mathrm{LLM}}(\bm{z}_{t}|\bm{y}_{<t})\!=\!\mathcal{N}(\phi_{\mathrm{LLM}}(\bm{y}_{<t}),\sigma^{2}\bm{I})$. The KL-divergence $\mathbb{D}_{\mathrm{KL}}(\mathbb{Q}_{U,\sigma^{2}}|\mathbb{P})$ is given by $\mathbb{D}_{\mathrm{KL}}(\mathbb{Q}_{U,\sigma^{2}}|\mathbb{P})\!=\!1/2\sigma^{2}\sum_{t=1}^{\tau}\!\bm{u}^{\top}_{t}\bm{u}_{t}$. After the optimization, we can obtain the optimized response $\bm{\mathrm{y}}_{\mathrm{AISP}}$ as $\bm{\mathrm{y}}_{\mathrm{AISP}}\!=\!\bm{\mathrm{y}}(V\!=\!U^{*})$ where $U^{*}$ is the solution of Eq. (3.1). Since we can generate several candidate responses at the last, we select $\bm{\mathrm{y}}_{\mathrm{AISP}}\!=\!\mathrm{arg}\!\max_{V\in\mathcal{V}}\bm{\mathrm{y}}(V)$ where $\mathcal{V}\!=\!\{V^{i}|V^{i}\!\sim\!q(V|U^{*},\sigma^{2})\}$ instead of just using $\bm{\mathrm{y}}(V\!=\!U^{*})$.
@@ 3.2 Free energy and Optimal distribution
Optimization problems such as Eq. (3.1) correspond to the optimal control problems called model predictive path integral control (MPPI) (Williams et al., 2018, 2017) and can be solved by considering a certain free energy. To optimize Eq. (3.1), we consider the following free energy:
@@
By using Jensen’s inequality, we have the following result:
@@
Free energy in Eq. (9) satisfies $\lambda F(r,p,\bm{\mathrm{x}},\lambda)\leq J(\bm{\mathrm{x}},U)$ and the equality holds if
@@
All of the proofs can be found in Appendix A. This theorem shows that the free energy in Eq. (9) is the lower bound of the objective function in Eq. (3.1), and the optimal density function is given by Eq. (10). The optimal density function in Eq. (10) is intractable, and it is difficult to obtain directly. Next, we show how to approximate it by using importance sampling. Note that while the similar result (Eq. (3)) has been presented for the distribution of responses, Eq. (10) is related to the distribution of pre-logits. This difference makes the optimization easy because $p(V)=p(V|\bm{0},\sigma)$ in Eq. (10) is a tractable distribution, and its sampling is easy as written in the next subsection.
@@ 3.3 Adaptive importance sampling
We consider to approximate the optimal distribution of Eq. (10) by the Gaussian distribution of Eq. (6) through importance sampling (Robert et al., 1999; Kloek and Van Dijk, 1978; Cappé et al., 2004; Bugallo et al., 2017). Let $\mathbb{Q}^{*}$ be the distribution corresponding to $q^{*}(V)$. We recall the following theorem, which was established in (Williams et al., 2018):
@@
$\textstyle w(V)\!=\!\frac{1}{\eta}\mathrm{exp}\!\left(\frac{1}{\lambda}r(\bm{\mathrm{x}},\bm{\mathrm{y}}(V))\!-\!\frac{1}{\sigma^{2}}\!\sum_{t=1}^{\tau}\hat{\bm{u}}_{t}^{\top}\bm{v}_{t}\!-\!\frac{1}{2\sigma^{2}}\hat{\bm{u}}_{t}^{\top}\hat{\bm{u}}_{t}\!\right)\!.$ (12)
@@
Note that our formulation is followed by MPPI (Williams et al., 2018) and can actually be extended to MPPI on pre-logits by changing sampling methods. While MPPI moves prediction and control windows by determining and applying $\hat{\bm{u}}_{1}$ to the control target for each iteration, AISP uses a fixed control window $t\!\in\![1,\tau]$. This is because once moving windows fix the prefix tokens, generated responses lose diversity. AISP can explore large response spaces by using the fixed window starting t=1t\!=\!1 and adaptive importance sampling.